# Masked Auto-Regressive Variational Acceleration: Fast Inference Makes Practical Reinforcement Learning

### v1

@@ 2.1 Hyperparameter Settings
The architecture of our one-step generator follows the same configuration of the original MAR models [zhou2024score]. The detailed specifications for the Denoising MLP used in the diffusion component are listed in Table 1.
@@
During the RL refinement stage, we use PickScore [kirstain2023pick] as the reward model. We generate samples with AR loops($N_{AR}$)=64 to calculate rewards, ensuring the optimization aligns with the final inference quality and requires less memory costs. Our best results in FID and IS are tested when inferring with $N_{AR}=128$.
@@ 3 Additional Results
This section provides additional qualitative results for the MARVAL-RL-L and MARVAL-RL-H models. These results were not included in the main paper for the following reasons:

### v2

@@ 2.1 Hyperparameter Settings
The architecture of our one-step generator follows the same configuration of the original MAR models [68].
@@
During the RL refinement stage, we use PickScore [22] as the reward model. We generate samples with AR loops($N_{AR}$)=64 to calculate rewards, ensuring the optimization aligns with the final inference quality and requires less memory costs. Our best results in FID and IS are tested when inferring with $N_{AR}=128$.
@@ 3 Additional Results
This section provides additional qualitative results for the MARVAL-RL-L and MARVAL-RL-H models, as well as extended text-to-image generation samples from our 1-step+RL DC-AR model. The MARVAL results were not included in the main paper for the following reasons:
@@
Furthermore, to supplement the text-to-image experiments presented in the main text, we provide a broader set of qualitative samples generated by our 1-step+RL DC-AR model in Fig. 3.
@@
Cheetah, pastel, cinematic lighting, full body view, stylize, detailed, v4.
The face of god, light, holy.
Fluffy koala bear happy.
Mechanical hummingbird made of brass and copper, hovering near flowers, intricate gears, steampunk style.
A watercolor of a beautiful forest with flowers, sunny, colorful.
A haunted house on a hill under a full moon.
Professional portrait of an anthropomorphic cat wearing gentleman hat and jacket walking in autumn forest.
Boy’s room, children’s room, comfortable room with toys, Freelancer game, drawing for a children’s storybook, blue room, cartoon.
Starfield, deep sky, realistic, 8k
Dragon made of constellation stars flying across night sky, over mountain landscape.
Skeleton, in the snow, friendly, made with pen and ink and prismacolor shading, blue and white color palet.
An elegant gold and white tikka head jewelry.
A framed renaissance painting of a unicorn on another planet in outer space. oil painting on canvass. rembrandt. 1647.
Christina Hendricks as a Disney princess, Full Torso, standing on a city street, octane render, volumetric lighting.
A painting of a rooster on a white background, a watercolor painting by Sun Kehong, shutterstock contest winner, cloisonnism, detailed painting, behance hd, photoillustration, 8K, UHD.
Starfield, space, 8k, photo, realism, sharp photography, maximum detail, sharp focus, Intricate details, epic, wide shot, highly realistic, cinematic lighting, volumetric lighting, octane render.
Figure 3: More qualitative results of distilled+RL DC-AR text-to-image generation. Our method achieves high generation quality and strong text alignment, demonstrating robust adaptability to prompts of varying lengths.